# The Nondecreasing Rank

### v1

@@ 6 Reduction of the ND rank to Nonnegative Rank
For invertible linear maps ${\mathbf{A}}_{j}$, we define the tensor product of these linear maps as the linear map ${\mathbf{A}}_{1}\otimes\cdots\otimes{\mathbf{A}}_{k}:\mathbb{R}^{p_{1}\times\cdots\times p_{k}}\rightarrow\mathbb{R}^{p_{1}\times\cdots\times p_{k}}$ that is given in coordinates by
$[({\mathbf{A}}^{(1)}\otimes\cdots\otimes{\mathbf{A}}^{(k)})({\mathbf{T}})]_{j_{1}\ldots j_{k}}=\sum_{i_{1}=1}^{p_{1}}\cdots\sum_{i_{k}=1}^{p_{k}}A_{j_{1}i_{1}}^{(1)}\cdots A_{j_{k}i_{k}}^{(k)}T_{i_{1}\ldots i_{k}}.$
On rank-one tensors this linear map is given by the simple formula $(\otimes_{j=1}^{k}{\mathbf{A}}^{(j)})(\otimes_{j=1}^{k}{\mathbf{v}}^{(j)})=\otimes_{j=1}^{k}{\mathbf{A}}^{(j)}{\mathbf{v}}^{(j)}$. The idea in this section is choose the maps ${\mathbf{A}}^{(j)}$ so that the order cone $\mathcal{C}(\mathcal{P}_{j})$ is mapped bijectively onto the positive orthant $\mathbb{R}^{p_{j}}_{+}$ by ${\mathbf{A}}^{(j)}$. As the tensor product map preserves rank-one tensors the ND rank and nonnegative rank will be equivalent whenever such ${\mathbf{A}}^{(j)}$ exist.
@@
By Lemma 2, $\mathcal{C}(\mathcal{P}_{j})$ is simplicial when $\mathcal{P}_{j}$ is a union of trees. One special case where the above theorem applies is when $\mathcal{P}=\times_{j=1}^{k}[p_{j}]$ under the standard ordering; the problem of finding an exact ND factorization can be converted into the problem of finding a nonnegative factorization. The matrices ${\mathbf{A}}^{(j)}$ required for this conversion are related to Möbius inversions. The Möbius tranform ${\mathbf{M}}_{\mathcal{P}}:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$ of a poset $\mathcal{P}$ is defined as the linear map with
@@
If $x\in\mathcal{S}_{i}$ then ${\mathbf{M}}_{\mathcal{P}}{\mathbf{e}}_{x}=\sum_{y\in\mathcal{P}:x\preceq y}{\mathbf{e}}_{y}=\sum_{y\in\mathcal{S}_{i}:x\preceq y}{\mathbf{e}}_{y}={\mathbf{M}}_{\mathcal{S}_{i}}{\mathbf{e}}_{x}$, where we have used the fact that the only elements in $\mathcal{P}$ that are comparable to $x$ are in $\mathcal{S}_{i}$. This proves the first statement. The second statement follows from
@@
where ${\mathbf{e}}_{(x_{1},\ldots,x_{k})}=\otimes_{j=1}^{k}{\mathbf{e}}_{x_{j}}$ is a rank-one, standard basis element of $\mathbb{R}^{\mathcal{P}}=\otimes_{j=1}^{k}\mathbb{R}^{\mathcal{P}_{j}}$. ∎
@@
When $\mathcal{P}=[p]$ is a chain the above formula simplifies to a Toeplitz matrix of the form:
@@
Putting these three results together, we see that if $\mathcal{P}_{j}=[p_{j}]$ is a chain then $\otimes_{j=1}^{k}{\mathbf{M}}_{\mathcal{P}_{j}}^{-1}$ maps $\mathcal{N}_{<\infty}$ onto $\otimes_{j=1}^{k}\mathbb{R}^{p_{j}}_{+}$, where an explicit formula for ${\mathbf{M}}_{\mathcal{P}_{j}}^{-1}$ is provided in (7). This provides a certificate of whether $\mathbf{T}$} is in $\mathcal{N}_{<\infty}$, since this occurs if and only if the entries of $(\otimes_{j=1}^{k}{\mathbf{M}}_{\mathcal{P}_{j}}^{-1})({\mathbf{T}})$ are nonnegative. The $(i_{1},\ldots,i_{k})$ entry of this tensor is
@@
where we use the convention that $T_{i_{1}-j_{1},\ldots,i_{k}-j_{k}}=0$ whenever $i_{l}-j_{l}=0$ for any $l$. Viewing $\mathbf{T}$} as a function, we can define the differencing operator along the $j$th mode as $(\Delta^{(j)}{\mathbf{T}})(i_{1},\ldots,i_{k})=T(i_{1},\ldots,i_{j},\ldots,i_{k})-T(i_{1},\ldots,i_{j}-1,\ldots,i_{k})$. From (8) the tensor $(\otimes_{j=1}^{k}\mathcal{M}_{\mathcal{P}_{j}}^{-1})({\mathbf{T}})$ can be seen to be equal to the tensor $\Delta^{(1)}\cdots\Delta^{(k)}{\mathbf{T}}$. In the matrix setting, with $p_{1}=2$, $p_{2}=3$ we obtain
$(\otimes_{j=1}^{2}{\mathbf{M}}_{\mathcal{P}_{j}}^{-1})({\mathbf{T}})$ $=\Delta^{(1)}\Delta^{(2)}\bigg{(}\begin{bmatrix}t_{11}&t_{12}&t_{13}\\ t_{21}&t_{22}&t_{23}\end{bmatrix}\bigg{)}=\Delta^{(1)}\bigg{(}\begin{bmatrix}t_{11}&t_{12}-t_{11}&t_{13}-t_{12}\\ t_{21}&t_{22}-t_{21}&t_{23}-t_{22}\end{bmatrix}\bigg{)}$
@@
Equations of the above form as were highlighted in [44], where monotonicity for functions defined on a Cartesian product of intervals were examined. The notion of a ${\mathbf{1}}_{k}$ increasing tensor in [44] is equivalent to the present notion of a finite ND rank tensor. Notice that similar expressions appeared in the previous section for the $\mathcal{H}$-representation for a product of collider posets.
We now discuss a probabilistic interpretation of the Möbius transform of the poset $\mathcal{P}=\times_{j=1}^{k}[p_{j}]$ and Theorem 4. Suppose that $\mathbf{R}$} is a nonnegative tensor with entries summing to one. The entry $R(i_{1},\ldots,i_{k})$ can be interpreted as a probability, where $\mathbf{R}$} a probability mass function (PMF) on $\mathcal{P}=\times_{j=1}^{k}[p_{j}]$. Any rank-$r$ nonnegative tensor factorization of $\mathbf{R}$} can be written in the form $\sum_{i=1}^{r}\lambda_{i}\otimes_{j=1}^{k}{\mathbf{q}}^{(ij)}$ with vectors $\bm{\lambda}\in\Delta_{r-1}$ and ${\mathbf{q}}^{(ij)}\in\Delta_{p_{j}-1}$ all residing in probability simplices. This means that $\mathbf{R}$} is a mixture model of $r$ probability mass functions $\otimes_{j=1}^{k}{\mathbf{q}}^{(ij)}$ on $\mathcal{P}$, with respective mixture weights $\lambda_{i}$ [21, Ch 4]. The tensor ${\mathbf{M}}_{\mathcal{P}}({\mathbf{R}})$ is exactly the multivariate conditional distribution function (CDF) of $\mathbf{R}$}, meaning that if ${\mathbf{X}}\in\mathcal{P}$ is a random variable with distribution given by $\mathbf{R}$} then ${\mathbf{M}}_{\mathcal{P}}({\mathbf{R}})(i_{1},\ldots,i_{k})=\text{Pr}({\mathbf{X}}\preceq(i_{1},\ldots,i_{k}))$. If ${\mathbf{F}}_{i}\in\mathcal{N}_{1}$ is the CDF corresponding to $\otimes_{j=1}^{k}{\mathbf{q}}^{(ij)}$ then ${\mathbf{M}}_{\mathcal{P}}({\mathbf{R}})=\sum_{i=1}^{r}\lambda_{i}{\mathbf{F}}_{i}$ has an ND rank of $r$. In summary, a probability distribution has a PMF with a nonnegative rank of $r$ if and only if it has a CDF with a ND rank of $r$.
To conclude this section we provide the $\mathcal{H}$-representation for $\mathcal{N}_{<\infty}$ when all but one cone $\mathcal{C}(\mathcal{P}_{j})$ is simplicial.
@@
$\sum_{i_{1}=x_{1},y_{1}}\cdots\sum_{i_{k}=x_{k},y_{k}}(-1)^{\sum_{j=1}^{k}\delta(i_{j})}T(i_{1},\ldots,i_{k})\geq 0,$ (9)
@@
By Lemma 1 $\big{(}\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})\big{)}^{*}=\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})^{*}$. The facets of $\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})$ are supported by hyperplanes corresponding to the extremal rays of $\big{(}\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})\big{)}^{*}$. Lemma 1 shows that the extremal rays of $\big{(}\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})\big{)}^{*}$ equal $\otimes_{j=1}^{k}{\mathbf{h}}^{(j)}$ where each ${\mathbf{h}}^{(j)}$ is an extremal ray of $\mathcal{C}(\mathcal{P}_{j})^{*}$. The set $\mathcal{C}(\mathcal{P}_{j})$ is the intersection of the halfspaces of the form $H_{xy}=\{{\mathbf{f}}:f_{x}\leq f_{y}\}$ for $x\preceq y$ and $x,y\in\mathcal{P}_{j}^{\prime}$. In the case where $x=0$ we define $H_{0y}=\{{\mathbf{f}}:0\leq f_{y}\}$. The intersection $H_{xy}\cap\mathcal{C}(\mathcal{P}_{j})$ is only $(p_{j}-1)$-dimensional when $x\lessdot y$ in $\mathcal{P}_{j}^{\prime}$, as if $x\prec z\prec y$ then $f_{x}=f_{z}=f_{y}$ for all $\mathbf{f}$} in $H_{xy}$ (where we have also defined $f_{0}\coloneq 0$) and $H_{xy}\cap\mathcal{C}(\mathcal{P}_{j})$ has dimension at most $p_{j}-2$. Conversely, if $x\lessdot y$ then $H_{ab}\cap\mathcal{C}(\mathcal{P}_{j})$ has dimension $p_{j}-1$.
@@
$\bigg{\{}{\mathbf{T}}:\otimes_{j=1}^{k}({\mathbf{e}}_{y_{j}}^{*}-{\mathbf{e}}_{x_{j}}^{*})({\mathbf{T}})=\sum_{i_{1}=x_{1},y_{1}}\cdots\sum_{i_{k}=x_{k},y_{k}}(-1)^{\sum_{j=1}^{k}\delta(i_{j})}T(i_{1},\ldots,i_{k})\geq 0\bigg{\}}.$
From this theorem we see that in the selenium example from Section 4, in addition to the monotonicity constraint $t_{12}\leq t_{32}$ that is violated in Table 1, the constraints $t_{11}-t_{12}-t_{31}+t_{32}\geq 0,\;\;t_{13}-t_{14}-t_{33}+t_{34}\geq 0$ are also violated; this matrix does not possess an exact ND decomposition.
@@ 7.1 The Maximum ND Rank
To find a lower bound on the maximum ND rank we use an analogue to the nested cone condition [28, Sec 2.1.1] for nonnegative matrix factorizations.
@@
The maximum ND rank of a matrix ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ is at least equal to the minimum number of elements ${\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{s}\in\mathcal{C}(\mathcal{P}_{1})$ needed for ${\mathbf{T}}(\mathcal{C}(\mathcal{P}_{2})^{*})\subseteq\mathrm{cone}({\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{s})$.
@@
If $\mathcal{C}(\mathcal{P}_{j})$ has $q_{j}$ extremal rays, $j=1,2$, and $\mathcal{C}(\mathcal{P}_{2})$ is simplicial then the maximum ND rank is $\min(q_{1},q_{2})$.
@@
The prototypical example of a poset with a non-simplicial order cone is a collider (Definition 4). The next theorem shows that when the column and row posets of a matrix both have the form of a collider the maximum ND rank can be significantly larger than even $\max(p_{1},p_{2})$.
@@
Let $\mathcal{P}=\{x_{1},\ldots,x_{p}\}$ where $x_{i}\prec x_{p}$ for all $i<p$. The maximum rank in $\mathcal{C}(\mathcal{P})\otimes\mathcal{C}(\mathcal{P})\subset\mathbb{R}^{p\times p}$ is $2^{p-1}$.
@@
Let $p\geq 3$ so that $\mathcal{P}$ contains a collider. The extremal rays ${\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{2^{p-1}}$, in $\mathcal{C}(\mathcal{P})$ can be taken to have the form $({\mathbf{w}},1)$ where $\mathbf{w}$} is any one of the $2^{p-1}$ vectors that have entries equal to either zero or one. Hence, $\mathcal{C}(\mathcal{P})$ is equal to the homogenization [56, Sec 1.5] of a hypercube. The facets of $\mathcal{C}(\mathcal{P})$ each contain $2^{p-2}$ extremal rays and there are 2(p1)2(p-1) facets; one facet for each equality $0=f_{x_{i}}$, and one facet for $f_{x_{i}}=f_{x_{p}}$, where $i<p$. The facet $\{{\mathbf{f}}:f_{x_{i}}=0\}\cap\mathcal{C}(\mathcal{P})$ contains the extremal rays of the form $({\mathbf{w}},1)$ with ${\mathbf{w}}_{i}=0$, while the facet $\{{\mathbf{f}}:f_{x_{i}}=f_{x_{p}}\}\cap\mathcal{C}(\mathcal{P})$ contains the remaining extremal rays $({\mathbf{w}},1)$ with $w_{i}=1$. These two facets are opposite to each other and partition the set of extremal rays.

### v2

@@ 6 Reduction of the ND rank to Nonnegative Rank
For invertible linear maps ${\mathbf{A}}^{(j)}$, we define the tensor product of these linear maps as the linear map ${\mathbf{A}}^{(1)}\otimes\cdots\otimes{\mathbf{A}}^{(k)}:\mathbb{R}^{p_{1}\times\cdots\times p_{k}}\rightarrow\mathbb{R}^{p_{1}\times\cdots\times p_{k}}$ that is given in coordinates by
$[({\mathbf{A}}^{(1)}\otimes\cdots\otimes{\mathbf{A}}^{(k)})({\mathbf{T}})]_{j_{1}\ldots j_{k}}=\sum_{i_{1}=1}^{p_{1}}\cdots\sum_{i_{k}=1}^{p_{k}}A_{j_{1}i_{1}}^{(1)}\cdots A_{j_{k}i_{k}}^{(k)}T_{i_{1}\ldots i_{k}}.$ (6)
On rank-one tensors this linear map is given by the simple formula $(\otimes_{j=1}^{k}{\mathbf{A}}^{(j)})(\otimes_{j=1}^{k}{\mathbf{v}}^{(j)})=\otimes_{j=1}^{k}{\mathbf{A}}^{(j)}{\mathbf{v}}^{(j)}$. The idea in this section is to choose the maps ${\mathbf{A}}^{(j)}$ so that the order cone $\mathcal{C}(\mathcal{P}_{j})$ is mapped bijectively onto the positive orthant $\mathbb{R}^{p_{j}}_{+}$ by ${\mathbf{A}}^{(j)}$. As the tensor product map preserves rank-one tensors, the problem of determining the ND rank can be transformed into a problem of determining the nonnegative rank whenever such ${\mathbf{A}}^{(j)}$ exist.
@@
By Lemma 2, $\mathcal{C}(\mathcal{P}_{j})$ is simplicial when the Hasse diagram of $\mathcal{P}_{j}$ is a union of trees. One special case where the above theorem applies is when $\mathcal{P}=\times_{j=1}^{k}[p_{j}]$ under the standard ordering; the problem of finding an exact ND factorization can be converted into the problem of finding a nonnegative factorization. The matrices ${\mathbf{A}}^{(j)}$ required for this conversion are related to Möbius inversions. The Möbius tranform ${\mathbf{M}}_{\mathcal{P}}:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$ of a poset $\mathcal{P}$ is defined as the linear map with
@@
We first prove that ${\mathbf{M}}_{\mathcal{P}}=\oplus_{i=1}^{m}{\mathbf{M}}_{\mathcal{S}_{i}}$, meaning that ${\mathbf{M}}_{\mathcal{P}}$ is a block-diagonal matrix with $m$ diagonal blocks that equal the ${\mathbf{M}}_{\mathcal{S}_{i}}$s. If $x\in\mathcal{S}_{i}$ then ${\mathbf{M}}_{\mathcal{P}}{\mathbf{e}}_{x}=\sum_{y\in\mathcal{P}:x\preceq y}{\mathbf{e}}_{y}=\sum_{y\in\mathcal{S}_{i}:x\preceq y}{\mathbf{e}}_{y}={\mathbf{M}}_{\mathcal{S}_{i}}{\mathbf{e}}_{x}$, where we have used the fact that the only elements in $\mathcal{P}$ that are comparable to $x$ are in $\mathcal{S}_{i}$. This proves the first statement. The second statement follows from
@@
where ${\mathbf{e}}_{(x_{1},\ldots,x_{k})}=\otimes_{j=1}^{k}{\mathbf{e}}_{x_{j}}$ is a rank-one, standard basis element of $\mathbb{R}^{\mathcal{P}}=\otimes_{j=1}^{k}\mathbb{R}^{\mathcal{P}_{j}}$. That ${\mathbf{M}}_{\mathcal{P}}=\otimes_{j=1}^{k}{\mathbf{M}}_{\mathcal{P}_{j}}$ is a consequence of the tensor product definition for linear maps provided in (6). ∎
@@
When $\mathcal{P}=[p]=\{1,\ldots,p\}$ is a chain the above formula simplifies to a Toeplitz matrix of the form:
@@
When each $\mathcal{P}_{j}=[p_{j}]=\{1,\ldots,p_{j}\}$ is a chain, Lemma 4 implies that ${\mathbf{M}}_{\mathcal{P}_{j}}^{-1}$ bijectively maps $\mathcal{C}(\mathcal{P}_{j})$ onto $\mathbb{R}^{p_{j}}_{+}$, while Lemma 3 further implies that $\otimes_{j=1}^{k}{\mathbf{M}}_{\mathcal{P}_{j}}^{-1}$ maps $\mathcal{N}_{<\infty}$ bijectively onto $\otimes_{j=1}^{k}\mathbb{R}^{p_{j}}_{+}$. An explicit formula for ${\mathbf{M}}_{\mathcal{P}_{j}}^{-1}$ is provided in (7). This provides a certificate of whether $\mathbf{T}$} is in $\mathcal{N}_{<\infty}$, since this occurs if and only if the entries of $(\otimes_{j=1}^{k}{\mathbf{M}}_{\mathcal{P}_{j}}^{-1})({\mathbf{T}})$ are nonnegative. The $(i_{1},\ldots,i_{k})$ entry of this tensor is
@@
where we use the convention that $T_{i_{1}-j_{1},\ldots,i_{k}-j_{k}}=0$ whenever $i_{l}-j_{l}=0$ for any $l$. Viewing $\mathbf{T}$} as a function, we can define the differencing operator along the $j$th mode as $(\Delta^{(j)}{\mathbf{T}})(i_{1},\ldots,i_{k})=T(i_{1},\ldots,i_{j},\ldots,i_{k})-T(i_{1},\ldots,i_{j}-1,\ldots,i_{k})$. From (8) the tensor $(\otimes_{j=1}^{k}{\mathbf{M}}_{\mathcal{P}_{j}}^{-1})({\mathbf{T}})$ can be seen to be equal to the tensor $\Delta^{(1)}\cdots\Delta^{(k)}{\mathbf{T}}$. In the matrix setting, with $p_{1}=2$, $p_{2}=3$ we obtain
[…]
@@
Equations of the above form were highlighted in [45], where monotonicity for functions defined on a Cartesian product of intervals were examined. The notion of a ${\mathbf{1}}_{k}$ increasing tensor in [45] is equivalent to the present notion of a finite ND rank tensor. Notice that similar expressions appeared in the previous section for the $\mathcal{H}$-representation for a product of collider posets.
We now discuss a probabilistic interpretation of the Möbius transform of the poset $\mathcal{P}=\times_{j=1}^{k}[p_{j}]$ and Theorem 4. Suppose that $\mathbf{R}$} is a nonnegative tensor with entries summing to one. The entry $R_{i_{1}\ldots i_{k}}$ can be interpreted as a probability, where $\mathbf{R}$} is a probability mass function (PMF) on $\mathcal{P}=\times_{j=1}^{k}[p_{j}]$. Any rank-$r$ nonnegative tensor factorization of $\mathbf{R}$} can be written in the form $\sum_{i=1}^{r}\lambda_{i}\otimes_{j=1}^{k}{\mathbf{q}}^{(ij)}$ with vectors $\bm{\lambda}\in\Delta_{r-1}$ and ${\mathbf{q}}^{(ij)}\in\Delta_{p_{j}-1}$ all residing in probability simplices. This means that $\mathbf{R}$} is a mixture of $r$ probability mass functions $\otimes_{j=1}^{k}{\mathbf{q}}^{(ij)}$ on $\mathcal{P}$, with respective mixture weights $\lambda_{i}$ [22, Ch 4]. The tensor ${\mathbf{M}}_{\mathcal{P}}({\mathbf{R}})$ is exactly the multivariate conditional distribution function (CDF) of $\mathbf{R}$}, meaning that if ${\mathbf{X}}\in\mathcal{P}$ is a random variable with distribution given by $\mathbf{R}$} then ${\mathbf{M}}_{\mathcal{P}}({\mathbf{R}})_{i_{1}\ldots i_{k}}=\text{Pr}({\mathbf{X}}\preceq(i_{1},\ldots,i_{k}))$. If ${\mathbf{F}}_{i}\in\mathcal{N}_{1}$ is the CDF corresponding to $\otimes_{j=1}^{k}{\mathbf{q}}^{(ij)}$ then ${\mathbf{M}}_{\mathcal{P}}({\mathbf{R}})=\sum_{i=1}^{r}\lambda_{i}{\mathbf{F}}_{i}$ has an ND rank of $r$. In summary, a probability distribution has a PMF with a nonnegative rank of $r$ if and only if it has a CDF with a ND rank of $r$.
To conclude this section we provide the $\mathcal{H}$-representation for $\mathcal{N}_{<\infty}$ when all but one cone $\mathcal{C}(\mathcal{P}_{j})$ is simplicial. The following theorem generalizes equation (8) beyond the case where each $\mathcal{P}_{j}$ is a chain. Furthermore, the left-hand side of equation (9) is the formula for obtaining a PMF from the CDF $\mathbf{T}$} when each $\mathcal{P}_{j}$ is a chain.
@@
$\sum_{i_{1}\in\{x_{1},y_{1}\}}\cdots\sum_{i_{k}\in\{x_{k},y_{k}\}}(-1)^{\sum_{j=1}^{k}\delta(i_{j})}T_{i_{1}\ldots i_{k}}\geq 0,$ (9)
@@
By Theorem 1 $\big(\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})\big)^{*}=\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})^{*}$. The facets of $\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})$ are supported by hyperplanes corresponding to the extremal rays of $\big(\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})\big)^{*}$. Lemma 1 shows that the extremal rays of $\big(\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})\big)^{*}$ equal $\otimes_{j=1}^{k}{\mathbf{h}}^{(j)}$ where each ${\mathbf{h}}^{(j)}$ is an extremal ray of $\mathcal{C}(\mathcal{P}_{j})^{*}$. The set $\mathcal{C}(\mathcal{P}_{j})$ is the intersection of the halfspaces of the form $\tilde{H}_{xy}=\{{\mathbf{f}}:f_{x}\leq f_{y}\}$ for $x\preceq y$ and $x,y\in\mathcal{P}_{j}^{\prime}$. In the case where $x=0$ we define $\tilde{H}_{0y}=\{{\mathbf{f}}:0\leq f_{y}\}$ and we let $H_{xy}$ be the hyperplane associated with the halfspace $\tilde{H}_{xy}$. The intersection $H_{xy}\cap\mathcal{C}(\mathcal{P}_{j})$ is only $(p_{j}-1)$-dimensional when $x\lessdot y$ in $\mathcal{P}_{j}^{\prime}$, as if $x\prec z\prec y$ then $f_{x}=f_{z}=f_{y}$ for all $\mathbf{f}$} in $H_{xy}$ (where we have also defined $f_{0}\coloneq 0$) and $H_{xy}\cap\mathcal{C}(\mathcal{P}_{j})$ has dimension less than $p_{j}-1$. Conversely, if $x\lessdot y$ then $H_{ab}\cap\mathcal{C}(\mathcal{P}_{j})$ has dimension $p_{j}-1$.
@@
$\bigg\{{\mathbf{T}}:\otimes_{j=1}^{k}({\mathbf{e}}_{y_{j}}^{*}-{\mathbf{e}}_{x_{j}}^{*})({\mathbf{T}})=\sum_{i_{1}\in\{x_{1},y_{1}\}}\cdots\sum_{i_{k}\in\{x_{k},y_{k}\}}(-1)^{\sum_{j=1}^{k}\delta(i_{j})}T_{i_{1}\ldots i_{k}}\geq 0\bigg\}.$
From this theorem we see that the matrix appearing in Table 1 in the selenium example does not possess an exact ND decomposition. This is because the constraints, $t_{12}\leq t_{32}$, $t_{11}-t_{12}-t_{31}+t_{32}\geq 0$, and $t_{13}-t_{14}-t_{33}+t_{34}\geq 0$, all of which are necessary for a ND factorization to exist, do not hold.
@@ 7.1 The Maximum ND Rank
If there are a large number of extremal rays in the given order cones the following lemma can provide a tighter upper bound on the maximum ND rank.
@@
The maximum ND rank is at most $\prod_{j=1}^{k}p_{j}$, where $\mathcal{C}(\mathcal{P}_{j})\subset\mathbb{R}^{p_{j}}$.
Proof.
By Caratheodory’s theorem, for any tensor ${\mathbf{T}}\in\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})=\text{cone}(\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}:{\mathbf{v}}^{(j)}\in\mathcal{C}(\mathcal{P}_{j}))$ the number of elements in $\{\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}:{\mathbf{v}}^{(j)}\in\mathcal{C}(\mathcal{P}_{j})\}$ needed to write $\mathbf{T}$} as a conic combination is at most $\prod_{j=1}^{k}p_{j}$. ∎
To find a lower bound on the maximum ND rank we use an analogue to the nested cone condition [29, Sec 2.1.1] for nonnegative matrix factorizations.
Lemma 8.
The ND rank of a matrix ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ is at least equal to the minimum number of elements ${\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{s}\in\mathcal{C}(\mathcal{P}_{1})$ needed for ${\mathbf{T}}(\mathcal{C}(\mathcal{P}_{2})^{*})\subseteq\mathrm{cone}({\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{s})$.
@@
If $\mathcal{C}(\mathcal{P}_{j})$ has $q_{j}$ extremal rays, $j=1,2$, and $\mathcal{C}(\mathcal{P}_{2})$ is simplicial then the maximum ND rank of a matrix in $\mathcal{N}_{<\infty}$ is $\min(q_{1},q_{2})$.
@@
The prototypical example of a poset with a non-simplicial order cone is a collider (Definition 4). The next theorem shows that when the column and row posets of a matrix both have the form of a collider the maximum ND rank can be larger than even $\max(p_{1},p_{2})$.
@@
Let $\mathcal{P}=\{x_{1},\ldots,x_{p}\}$ where $x_{i}\prec x_{p}$ for all $i<p$. The maximum rank in $\mathcal{C}(\mathcal{P})\otimes\mathcal{C}(\mathcal{P})\subset\mathbb{R}^{p\times p}$ is at least 2(p1)2(p-1). When $p=3$ the maximum rank is equal to $4$.
@@
Let $p\geq 3$ so that $\mathcal{P}$ contains a collider. The extremal rays ${\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{2^{p-1}}$, in $\mathcal{C}(\mathcal{P})$ can be taken to have the form $({\mathbf{w}},1)$ where ${\mathbf{w}}\in\mathbb{R}^{p-1}$ is any one of the $2^{p-1}$ vectors that have entries equal to either zero or one. Hence, $\mathcal{C}(\mathcal{P})$ is equal to the homogenization [59, Sec 1.5] of a hypercube. The facets of $\mathcal{C}(\mathcal{P})$ each contain $2^{p-2}$ extremal rays and there are 2(p1)2(p-1) facets; one facet for each equality $0=f_{x_{i}}$, and one facet for $f_{x_{i}}=f_{x_{p}}$, where $i<p$. The facet $\{{\mathbf{f}}:f_{x_{i}}=0\}\cap\mathcal{C}(\mathcal{P})$ contains the extremal rays of the form $({\mathbf{w}},1)$ with ${\mathbf{w}}_{i}=0$, while the facet $\{{\mathbf{f}}:f_{x_{i}}=f_{x_{p}}\}\cap\mathcal{C}(\mathcal{P})$ contains the remaining extremal rays $({\mathbf{w}},1)$ with $w_{i}=1$. These two facets are opposite to each other and partition the set of extremal rays.
The goal of the remainder of the proof to construct a matrix $\mathbf{T}$} that has an ND rank bounded below by 2(p1)2(p-1). Choose a facet $F$ of $\mathcal{C}(\mathcal{P})$ that has $F^{\prime}$ as its opposite facet. Let $J$ be an index set of size less than p1p-1 where ${\mathbf{v}}_{j}\in F$ for every $j\in J$. The set $\text{cone}({\mathbf{v}}_{j}:j\in J)$ has a dimension that is less than p1p-1. If we remove all such cones from $F$ we are left with a non-empty set
$F\;\backslash\bigg(\bigcup_{J:|J|<p-1,\;\forall j\in J\;{\mathbf{v}}_{j}\in F}\text{cone}({\mathbf{v}}_{j}:j\in J)\bigg),$
because $F$ has dimension p1p-1. Choose a point $\sum_{i\in I_{F}}\alpha_{i}{\mathbf{v}}_{i}$ in the above set, where the index subset has $|I_{F}|=p-1$ and ${\mathbf{v}}_{i}\in F$ for every $i\in I_{F}$. Repeat the above process for the facet $F^{\prime}$ to obtain an analogous point $\sum_{i\in I_{F^{\prime}}}\alpha_{i}{\mathbf{v}}_{i}$. We define the matrix ${\mathbf{T}}=\sum_{i\in I}\alpha_{i}{\mathbf{v}}_{i}\otimes{\mathbf{v}}_{i}$ where $I=I_{F}\cup I_{F^{\prime}}$.